OpenStudy (anonymous):

find the equation of the line having the given slope containing the given point m=-4, (3,9)

jimthompson5910 (jim_thompson5910):

\[\Large y = mx+b\] \[\Large 9 = \left(-4\right)\left(3\right)+b\] \[\Large 9 = -12+b\] \[\Large 9+12 = b\] \[\Large 21 = b\] \[\Large b = 21\] So the equation of the line is \[\Large y = -4x+21\]